# Starlight from JWST: Implications for star formation and dark matter models

### v2

@@ 5.1 Cold dark matter
The full posteriors for our CDM fit parameters are shown in Fig. 10 and the best-fit values are given in Table 1. Most of the parameters, in particular those parametrizing the broken power-law shape of $f_{*}(M)$, are well constrained by the data, and the posteriors show only a mild negative correlation between $\alpha$ and $\beta$ as well as between $\alpha$ and $M_{c}$. The fits of $\alpha$, $\beta$ and $M_{c}$ are similar to those found by Harikane et al. (2022) but we find about a factor of two higher SFR. Compared to Sabti et al. (2022a), we have fixed the cosmological parameters to the Planck CMB values and we are able to constrain $M_{c}$ and $\beta$ better because the HST data we consider extend to higher luminosities.
From the posteriors of $\gamma$, $f_{\kappa}$ and $z_{\kappa}$, we see a clear ($>95\%$) preference for a relatively rapid change in $\kappa_{\rm UV}$ around $z_{\kappa}=10-11$ by factor of $\sim 3$. The posterior of $z_{\rm fb}$ shows that the preference for a high-$z$ change in the powers $\alpha$ and $\beta$ is less significant, $<95\%$. This is expected, since the change is driven by the measurements at $z=17$ and $z=25$ that have relatively large uncertainties and $\Phi=0$ is within the 95% range. Moreover, we see a correlation between $f_{\kappa}$ and $z_{\rm bf}$, indicating that the enhancement of the UV luminosity can be partly reduced by changing the powers $\alpha$ and $\beta$. The data prefer a narrow scatter in the UV magnitude, $\sigma_{\rm UV}<0.15$ at the $95\%$ CL, and give an upper bound on the scale of the exponential SFR suppression, $M_{t}<10^{8.78}\,M_{\odot}$ at the $95\%$ CL. This implies that the explanation for the UV excess as a consequence of a large constant scatter in the $(M_{\rm UV}\,,M)$ is disfavoured when compared to the Pop-III star hypothesis.  [^6] However, a mass-dependent scatter that increases at small halo masses could enhance the UV luminosities at high-$z$ (Gelli et al. 2024). This result should be treated with caution, however, since selection effects, which are not considered in the likelihood analysis, could modify the conclusion.
The evolution of $\kappa_{\rm UV}$ for the best fit is shown in the upper left panel of Fig. 6. The range shown in gray corresponds to estimates of the conversion factor for Pop-III stars with top-heavy initial mass functions (Harikane et al. 2023). We see that the best fit prefers a sharp transition from Pop-I and Pop-II stars with standard Salpeter profiles up to $z\sim 10$ to a top-heavy Pop-III population at $z>10.8$. The upper right panel of Fig. 6 illustrates how the feedback effects are suppressed for the best fit. This strongly enhances the SFR density,
@@
In Fig. 7 we show the best-fit UV luminosity function in the CDM model (in black) compared with the HST and JWST observations at different redshifts. The change in the tail shape at $M_{\rm UV}<-22$ is primarily driven by the lensing magnification. The slight turn seen in the last panel at $M_{\rm UV}>-18$ is due to the exponential term in (34) that becomes relevant in the shown $M_{\rm UV}$ range only at the highest $z$, as seen from the lower right panel of Fig. 6. The fit without the suppression of the feedback effects is shown by the dashed lines, and undershoots the measurements at $z=17$ and $25$. The dot-dashed curve shows the fit when also the high-$z$ enhancement in the luminosities is removed. We see that this case undershoots the observed luminosity function at redshifts $z>10$.
@@ 5.2 Suppressed small scale perturbations
We also see in Fig. 7 the comparison between the UV data and the predictions with suppressed small-scale structures in the FDM (green) and WDM (orange) models. For FDM we show the case where the model parameters correspond to the CDM best fit and the FDM mass is at the 95% CL lower limit, while for WDM we show the best fit case which is slightly better than in the CDM model. We see that the predictions are indistinguishable for $z\lesssim 10$, and up to $z\sim 15$ the suppression is at lighter halos (bigger $M_{\rm UV}$) than what the observations can probe. At the highest redshifts, $z>15$, we see that the model predictions undershoot the observations. However, since the uncertainties in the data are large, the observations at high-$z$ are compatible with a strongly-suppressed UV luminosity function.
@@
For the FDM model, the parameters for the star formation rate, the enhancement of the UV luminosity and the suppression of the feedback effects remain the same as in the CDM model, as seen in the posterior plots in Fig. 11. This reflects the fact that the FDM prediction deviates sharply from the CDM prediction at the suppression scale. The main difference between FDM and CDM models is seen in $z_{*}$, which determines when the feedback effects are fully removed. For small FDM masses the feedback effects can be fully removed at lower $z$ because the luminosity function is suppressed by the FDM quantum pressure, which causes the peak in the posterior of the FDM mass shown in the left panel of Fig. 8. The posterior also shows a lower bound on the FDM mass, $m_{\rm FDM}>4.5\times 10^{-22}\,{\rm eV}$ at 95% CL. The UV luminosity function has been previously studied in the FDM model in (Bozek et al. 2015; Schive et al. 2016; Corasaniti et al. 2017; Winch et al. 2024). Our constraint on the FDM mass is somewhat stronger than that obtained by Winch et al. (2024) using the HST data and JWST data from Harikane et al. (2024).
@@
The deviation from the CDM prediction is smoother in the WDM case. From the posterior plots in Fig. 12, we see that the WDM model favours a narrow mass range around 2 keV and the fit at small $m_{\rm WDM}$ moves to smaller values of $\alpha$ and larger values of $\beta$ than in the CDM model. The maximal likelihood in the WDM model is slihgtly higher than in CDM, $\Delta\ln\mathcal{L}\approx 2.3$, but the Bayesian Information Criterion (BIC), which accounts for the increase in the number of parameters, favours CDM over WDM with $\Delta{\rm BIC}\approx 0.7$. The posterior of the WDM mass shown in the middle panel of Fig. 8 shows a lower bound $m_{\rm WDM}>1.5\,{\rm keV}$ at 95% CL. Several previous works have derived constraints on WDM from UV luminosity function observations (Menci et al. 2016; Corasaniti et al. 2017; Rudakovskyi et al. 2021; Hibbard et al. 2022; Maio & Viel 2023; Liu et al. 2024). Although we use a larger set of JWST data, the bound on the WDM mass is weaker than what was found by Liu et al. (2024). This difference arises because Liu et al. (2024) uses a low-$z$ fit of the HMF to numerical simulations derived in Stücker et al. (2021), while we estimate the WDM HMF using the transfer function and the smooth-$k$ window function.
@@ 5.3 Enhanced small-scale perturbations
Enhancing the small-scale structure has the opposite effect as WDM or FDM, but the conclusions are very similar. We see in Fig. 7 that for $z\lesssim 10$ the predictions resemble those of CDM, and only at the highest redshifts do we see a small enhancement of the UV luminosity function that growths at lower scales, $M_{\rm UV}\lesssim-20$. This reflects the fact that the higher the redshift, the more deviations from CDM one can expect to see, especially for low-mass halos. When fitting the star formation parameters, the results are independent of the enhancement of the scales and very similar to CDM, as seen in Fig. 13. We conclude that enhancing the matter power spectrum does not help to explain the excess in the UV radiation at high-$z$ when other possibilities like Pop-III stars are taken into account. The fit favors a sharp transition to Pop-III stars and such a feature cannot be reproduced by enhancing the matter power spectrum. The fit gives a lower bound $k_{c}>27\,{\rm Mpc}^{-1}$ at 95% CL, as seen in the right panel of Fig. 8. The excluded region corresponds to axion miniclusters produced with $m_{a}<7.5\times 10^{-17}\,{\rm eV}$ or PBHs that obey $f_{\rm PBH}>\max[88M_{\odot}/m_{\rm PBH},10^{-4}(m_{\rm PBH}/10^{4}M_{\odot})^{-0.09}]$ where the second bound comes from the requirement that $k_{c}<k_{\rm cut}$. Our constraint on $k_{c}$ is significantly stronger than that implied by the measurements of the matter power spectrum at scales $100.5<k/{\rm Mpc}^{-1}<$ 10 derived by Sabti et al. (2022b) and that implied by the analysis of Iršič et al. (2020), both using the HST data alone. As for PBHs, our constraint is stronger than that derived by Murgia et al. (2019), which considered the limit on an enhanced small-scale power spectrum imposed by the Lyman-$\alpha$ forest observations. In Fig. 9 we show a compilation of previous PBH constraints together with our new constraint in red.
[…]
Figure 9: Constraints on the fraction of the DM density provided by PBHs from the OGLE microlensing survey (Mróz et al. 2024), accretion limits from CMB observations (Serpico et al. 2020), SNe lensing (Zumalacarregui & Seljak 2018), the population of X-ray sources (Inoue & Kusenko 2017), Lyman-$\alpha$ data (Murgia et al. 2019), survival of wide binaries (Monroy-Rodríguez & Allen 2014), the population of gravitational wave (GW) sources (Andrés-Carcasona et al. 2024) and GW lensing (Urrutia et al. 2023). The red line shows the upper limit derived in this work from measurements of the UV luminosity function.
@@ 6 Conclusions
We have investigated in this work the implications of the high-$z$ UV luminosity function observations with HST and JWST for star formation and different DM models. We have revisited the excursion-set formalism to compute the unconditional and conditional halo mass functions and derived a new estimate of the halo growth rate. We have also proposed a new window function, inspired by the smooth-$k$ Leo et al. (2018) window function, that reproduces well the numerical results and matches better the CDM predictions at high masses.
We have found that, assuming that the SFR is proportional to the halo growth rate, the observations of the UV luminosity function indicate two interesting effects that change the luminosity function at high $z$. First, we find evidence at more than $95\%$ CL for a sharp enhancement in the luminosity function around $z\simeq 10$. This can be explained by a transition from a Salpeter mass function to a top-heavy Pop-III mass function. Beyond $z=10-15$ the luminosity function needs to be further enhanced, in particular to match the JWST observations at $z=17$ and 25. This enhancement can be achieved by reducing the feedback from SNe and AGNs. The evidence for this second effect is, however, less than $95\%$ CL. We have also considered the effect of dust attenuation and lensing, and shown that they are relevant for the bright end of the luminosity function at $z<10$ but do not play a significant role in fitting the higher redshift data.
We have considered departures from CDM in two directions, suppressing the small-scale structures via either warm or fuzzy DM and enhancing the matter power spectrum with white noise, which could happen in scenarios including axion miniclusters or PBHs. When marginalizing over the SFR parametrization, we find a lower bound of $4.5\times 10^{-22}{\rm eV}$ for the FDM mass and $1.5\,{\rm keV}$ for the WDM mass, at $95\%$ CL. Because the new JWST data at high-$z$ still have significant uncertainties, we find that the DM bounds are not much stronger than those obtained using the HST data alone. Another important conclusion is that enhanced matter perturbations are not preferred over changes in the star formation rate due to the presence of Pop-III stars and changes in the feedback in the pre-reionisation era. We have found that the data constrain the scale of the enhancement to $k_{c}>27\,{\rm Mpc}^{-1}$ at $95\%$ CL, excluding, for example, the possibility of axion miniclusters produced for $m_{a}<7.5\times 10^{-17}\,{\rm eV}$ or heavy PBHs that constitute a fraction $f_{\rm PBH}>\max[88M_{\odot}/m_{\rm PBH},10^{-4}(m_{\rm PBH}/10^{4}M_{\odot})^{-0.09}]$ of DM.
@@ Acknowledgements.
This work was supported by the Estonian Research Council grants PRG803, PSG869, RVTT3 and RVTT7 and the Center of Excellence program TK202. The work of J.E. and M.F. was supported by the United Kingdom STFC Grants ST/T000759/1 and ST/X000753/1. The work of V.V. was partially supported by the European Union’s Horizon Europe research and innovation program under the Marie Skłodowska-Curie grant agreement No. 101065736.
Appendix A Corner plots
The full posteriors of the MCMC scans over the model parameters in the four models considered in this paper are shown in Figs. 10, 11, 12 and 13.
Figure 10: Posteriors of SFR fit to the UV luminosity function observations in the CDM model. The contours indicate the 68% and 95% credible regions.
Figure 11: Posteriors of SFR fits to the UV luminosity function observations in the FDM model. The contours indicate the 68% and 95% credible regions.
Figure 12: Posteriors of SFR fits to the UV luminosity function observations in the WDM model. The contours indicate the 68% and 95% credible regions. The parameter $m$ corresponds to $m_{\rm FDM}/10^{-22}\,{\rm eV}$ in the FDM model and to $m_{\rm WDM}/{\rm keV}$ in the WDM model.
Figure 13: Posteriors of SFR fits to the UV luminosity function observations in models with a white noise-dominated matter power spectrum at $k>k_{c}$. The contours indicate the 68% and 95% credible regions.

### v4

@@ 5.1 Cold dark matter
The full posteriors for our CDM fit parameters are shown in blue in Fig. 7 and the best-fit values are given in Table 6. We have checked that the Gelman-Rubin statistic $R$ (Gelman:1992zz) is very close to $1$ for each of the parameters. This indicates good convergence of the MCMC chains. The highest value, $R\approx 1.09$, was obtained for $\log_{10}(M_{t}/M_{\odot})$.
Most of the parameters, in particular those parametrising the broken power-law shape of $f_{*}(M)$, are well constrained by the data, and the posteriors show only a mild negative correlation between $\alpha$ and $\beta$ as well as between $\alpha$ and $M_{c}$. The fits of $\alpha$, $\beta$ and $M_{c}$ are similar to those found by 2022ApJS..259...20H, but we find about a factor of two higher SFR. Compared to Sabti:2021xvh, we have fixed the cosmological parameters to the Planck CMB values and we are able to constrain $M_{c}$ and $\beta$ better because the HST data we consider extend to higher luminosities.
From the posteriors of $\gamma$, $f_{\kappa}$ and $z_{\kappa}$, we see a clear ($95\%$) preference for a relatively rapid change in $\kappa_{\rm UV}$ around $z_{\kappa}=10-11$ by factor of $\sim 3$. The posterior of $z_{\rm fb}$ shows that the preference for a high-$z$ change in the powers $\alpha$ and $\beta$ is less significant, $95\%$. This is expected, since the change is driven by the measurements at $z=17$ and $z=25$ that have relatively large uncertainties and $\Phi=0$ is within the 95% range. Moreover, we see a correlation between $f_{\kappa}$ and $z_{\rm bf}$, indicating that the enhancement of the UV luminosity can be partly reduced by changing the powers $\alpha$ and $\beta$. The data prefer a narrow scatter in the UV magnitude, $\sigma_{\rm UV}<0.15$ at the $95\%$ CL, and give an upper bound on the scale of the exponential SFR suppression, $M_{t}<10^{9}\,M_{\odot}$ at the $95\%$ CL. This implies that the explanation for the UV excess as a consequence of a large constant scatter in the $(M_{\rm UV}\,,M)$ is disfavoured when compared to the Pop-III star hypothesis.  [^7] However, a mass-dependent scatter that increases at small halo masses could enhance the UV luminosities at high-$z$ (2024ApJ...975..192G). This result should be treated with caution, however, since selection effects, which are not considered in the likelihood analysis, could modify the conclusion.
The evolution of $\kappa_{\rm UV}$ for the best fit is shown in the upper left panel of Fig. 6. The range shown in gray corresponds to estimates of the conversion factor for Pop-III stars with top-heavy initial mass functions (Harikane:2022rqt). We see that the best fit prefers a sharp transition from Pop-I and Pop-II stars with standard Salpeter profiles up to $z\sim 10$ to a top-heavy Pop-III population at $z>10.8$. The upper right panel of Fig. 6 illustrates how the feedback effects are suppressed for the best fit. This strongly enhances the SFR density,
@@
In Fig. 8 we show the best-fit UV luminosity function in the CDM model (in black) compared with the HST and JWST observations at different redshifts. The change in the tail shape at $M_{\rm UV}<-22$ is primarily driven by the lensing magnification. The slight turn seen in the last panel at $M_{\rm UV}>-18$ is due to the exponential term in (34) that becomes relevant in the shown $M_{\rm UV}$ range only at the highest $z$, as seen from the lower right panel of Fig. 6. The fit without the suppression of the feedback effects is shown by the dashed lines, and undershoots the measurements at $z=17$ and $25$. The dot-dashed curve shows the fit when also the high-$z$ enhancement in the luminosities is also removed. We see that this case undershoots the observed luminosity function at redshifts $z>10$. Moreover, we see that our model does not provide a good fit of the brightest objects at $z=17$. This may reflect a limitation of our parametrisation, but it is also possible that the luminosities of some of these objects are contaminated by AGNs (Castellano:2025vkt).
Figure 8: The UV luminosity function at different redshifts in the cold (black), fuzzy (green) and warm (orange) DM models and in the model with enhanced matter power (red). The best fit is shown for CDM and WDM cases, while for FDM the mass corresponds to the 95% lower bound. The dashed and dot-dashed curves show the CDM UV luminosity function without the suppression of the feedback effects and without the change in the UV conversion factor $\kappa_{\rm UV}$. The points with errorbars show the HST measurements from 2021AJ....162...47B (cyan) and 2022ApJS..259...20H (pink) and the JWST measurements from 2024MNRAS.533.3222D (blue), Perez-Gonzalez:2025bqr (brown) and Castellano:2025vkt (gray).
Figure 9: Posteriors of SFR fits to the UV luminosity function observations as a function of FDM mass (top panels), WDM mass (middle panels) and the enhancement scale (bottom panels). The contours indicate the 68% and 95% credible regions.
@@ 5.2 Suppressed small scale perturbations
We also see in Fig. 8 the comparison between the UV data and the predictions with suppressed small-scale structures in the FDM (green) and WDM (orange) models. For FDM we show the case where the model parameters correspond to the CDM best fit and the FDM mass is at the 95% CL lower limit, while for WDM we show the best fit case, which is slightly better than in the CDM model. We see that the predictions are indistinguishable for $z\lesssim 10$, and up to $z\sim 15$ the suppression is at lighter halos (bigger $M_{\rm UV}$) than what the observations can probe. At the highest redshifts, $z>15$, we see that the model predictions undershoot the observations. However, since the uncertainties in the data are large, the observations at high-$z$ are compatible with a strongly suppressed UV luminosity function.
@@
For the FDM model, the parameters for the star formation rate, the enhancement of the UV luminosity and the suppression of the feedback effects remain the same as in the CDM model, as seen in the posterior plots in Fig. 7. This reflects the fact that the FDM prediction deviates sharply from the CDM prediction at the suppression scale. The main difference between FDM and CDM models is seen in $z_{*}$, which determines when the feedback effects are fully removed. For small FDM masses the feedback effects can be fully removed at lower $z$ because the luminosity function is suppressed by the FDM quantum pressure, which causes the peak in the posterior of the FDM mass shown in the top panels of Fig. 9. The posterior also shows a lower bound on the FDM mass, $m_{\rm FDM}>5.6\times 10^{-22}\,{\rm eV}$ at 95% CL. The UV luminosity function has been previously studied in the FDM model in (Bozek:2014uqa; Schive:2015kza; Corasaniti:2016epp; Winch:2024mrt; Lazare:2024uvj). Our constraint on the FDM mass is somewhat stronger than that obtained by Winch:2024mrt using the HST data and JWST data from 2024ApJ...960...56H.
The deviation from the CDM prediction is smoother in the WDM case, and starts already much above the half-mode mass. Consequently, the WDM fit at small $m_{\rm WDM}$ moves to smaller values of $\alpha$ and larger values of $\beta$ than in the CDM model, as seen in the posterior plots in Figs. 7 and 9. Due to the degeneracy between $\alpha$ and $m_{\rm WDM}$, in particular, the WDM model favours a narrow mass range around 2 keV. The maximal likelihood in the WDM model is slightly higher than in CDM, $\Delta\ln\mathcal{L}\approx 2.7$, and the Bayesian Information Criterion (BIC), which accounts for the increase in the number of parameters, indicates a mild (insignificant) preference for WDM over CDM with $\Delta{\rm BIC}\approx-0.4$. As seen from Fig. 8, the fit improves in particular at $z=17$.
@@
The posteriors of the WDM mass shown in the middle panels of Fig. 9 show a lower bound $m_{\rm WDM}>1.5\,{\rm keV}$ at 95% CL. Several previous works have derived constraints on WDM from UV luminosity function observations (Menci:2016eui; Corasaniti:2016epp; Rudakovskyi:2021jyf; Hibbard:2022sng; Maio:2022lzg; Liu:2024edl). Although we use a larger set of JWST data, the bound on the WDM mass is weaker than what was found by Liu:2024edl. This difference arises because Liu:2024edl uses a low-$z$ fit of the HMF to numerical simulations derived in Stucker:2021vyx, while we estimate the WDM HMF using the transfer function and the smooth-$k$ window function.
@@ 5.3 Enhanced small-scale perturbations
Enhancing the small-scale structure has the opposite effect as WDM or FDM, but the conclusions are very similar. We see in Fig. 8 that for $z\lesssim 10$ the predictions resemble those of CDM, and only at the highest redshifts do we see a small enhancement of the UV luminosity function that grows at lower scales, $M_{\rm UV}\lesssim-20$. This reflects the fact that the higher the redshift, the more deviations from CDM one can expect to see, especially for low-mass halos. When fitting the star formation parameters, the results are independent of the enhancement of the scales and very similar to CDM, as seen in Fig. 7. We conclude that enhancing the matter power spectrum does not help to explain the excess in the UV radiation at high-$z$ when other possibilities like Pop-III stars are taken into account. The fit favours a sharp transition to Pop-III stars and such a feature cannot be reproduced by enhancing the matter power spectrum. The fit gives a lower bound $k_{c}>25\,{\rm Mpc}^{-1}$ at 95% CL, as seen in the bottom panels of Fig. 9. The excluded region corresponds to axion miniclusters produced with $m_{a}<6.6\times 10^{-17}\,{\rm eV}$ or PBHs that obey $f_{\rm PBH}>\max[105M_{\odot}/m_{\rm PBH},10^{-4}(m_{\rm PBH}/10^{4}M_{\odot})^{-0.09}]$ where the second bound comes from the requirement that $k_{c}<k_{\rm cut}$. Our constraint on $k_{c}$ is significantly stronger than that implied by the measurements of the matter power spectrum at scales $0.5<k/{\rm Mpc}^{-1}<10$ derived by Sabti:2021unj and that implied by the analysis of Irsic:2019iff, both using the HST data alone. As for PBHs, our constraint is stronger than that derived by Murgia:2019duy, which considered the limit on an enhanced small-scale power spectrum imposed by the Lyman-$\alpha$ forest observations, and that derived by Gouttenoire:2023nzr using the HST UV luminosity function data. In Fig. 10, we show a compilation of previous PBH constraints together with our new constraint in red.
Figure 10: Constraints on the fraction of the DM density provided by PBHs from the OGLE microlensing survey (Mroz:2024mse), accretion limits from CMB observations (Serpico:2020ehh), SNe lensing (Zumalacarregui:2017qqd), the population of X-ray sources (Inoue:2017csr), Lyman-$\alpha$ data (Murgia:2019duy), survival of wide binaries (Monroy-Rodriguez:2014ula), the population of gravitational wave (GW) sources (Andres-Carcasona:2024wqk) and GW lensing (Urrutia:2023mtk). The red line shows the upper limit derived in this work from measurements of the UV luminosity function.
@@ 6 Conclusions
We have investigated in this work the implications of the high-$z$ UV luminosity function observations with HST and JWST for star formation and different DM models. We have revisited the excursion-set formalism to compute the unconditional and conditional halo mass functions and derived a new estimate of the halo growth rate. We have also proposed a new window function, inspired by the smooth-$k$ Leo:2018odn window function, that reproduces well the numerical results and matches better the CDM predictions at high masses.
We have found that, assuming that the SFR is proportional to the halo growth rate, the observations of the UV luminosity function indicate two interesting effects that change the luminosity function at high $z$. First, we find evidence at more than $95\%$ CL for a sharp enhancement in the luminosity function around $z\simeq 10$. This can be explained by a transition from a Salpeter mass function to a top-heavy Pop-III mass function. Beyond z=1015z=10-15 the luminosity function needs to be further enhanced, in particular to match the JWST observations at $z=17$ and 25. This enhancement can be achieved by reducing the feedback from SNe and AGNs. The evidence for this second effect is, however, less than $95\%$ CL. We have also considered the effect of dust attenuation and lensing, and shown that they are relevant for the bright end of the luminosity function at $z<10$ but do not play a significant role in fitting the higher redshift data.
We have considered departures from CDM in two directions, suppressing the small-scale structures via either warm or fuzzy DM and enhancing the matter power spectrum with white noise, which could happen in scenarios including axion miniclusters or PBHs. When marginalising over the SFR parametrisation, we find a lower bound of $5.6\times 10^{-22}{\rm eV}$ for the FDM mass and $1.5\,{\rm keV}$ for the WDM mass, at $95\%$ CL. Because the new JWST data at high-$z$ still have significant uncertainties, we find that the DM bounds are not much stronger than those obtained using the HST data alone. Another important conclusion is that enhanced matter perturbations are not preferred over changes in the star formation rate due to the presence of Pop-III stars and changes in the feedback in the pre-reionisation era. We have found that the data constrain the scale of the enhancement to $k_{c}>25\,{\rm Mpc}^{-1}$ at $95\%$ CL, excluding, for example, the possibility of axion miniclusters produced for $m_{a}<6.6\times 10^{-17}\,{\rm eV}$ or heavy PBHs that constitute a fraction $f_{\rm PBH}>\max[105M_{\odot}/m_{\rm PBH},10^{-4}(m_{\rm PBH}/10^{4}M_{\odot})^{-0.09}]$ of DM.
@@ Acknowledgements.
This work was supported by the Estonian Research Council grants PRG803, PSG869, RVTT3 and RVTT7 and the Center of Excellence program TK202. The work of J.E. and M.F. was supported by the United Kingdom STFC Grants ST/T000759/1 and ST/X000753/1. The work of $V.$ was partially supported by the European Union’s Horizon Europe research and innovation program under the Marie Skłodowska-Curie grant agreement No. 101065736.